[Ncep.nhc.nco_contacts] TSB Morning Rounds - Thu Nov 10, 2016

Craig Mattocks - NOAA Federal craig.mattocks at noaa.gov
Thu Nov 10 16:11:50 UTC 2016


Here is today's summary of NHC computer operations:

--- Craig Mattocks

—————

TAFB:

1. Jeff Lewitsky noticed that our NCO internet connection was down at about 22:22 UTC (5:22 pm EST) yesterday. He reported this to the NOAA SDM. Service was restored about 20 minutes later.

2. The NOAA SDM called around 22:20 UTC (5:20 pm EST) to inquire if NHC was having any data problems. They were seeing alarms on NHC's line 162 socket feed. Matt Sardi was contacted and ran 'ldmadmin watch' on logger.nhc.noaa.gov - no data was coming in. He also stopped and started ldmadmin, but still did not see any products coming in. He recommended that Scott Stripling follow up with the SDM and ask them to have the TOC check upstream for data outages on their end.

Tech control called around 07:15 UTC to tell TAFB that line 162 was out again (trouble ticket #151109-0152). No obvious impacts were observed.

It turned out that these problems were caused by the Century Link circuit migration (where all the centers connect to NCEP), upgrade to 10G speeds, and replacement/reconfiguration of the network routers. This caused "asymmetric routing" between some of the centers, slowing down/diverting traffic southbound. Once the network was reverted to an old legacy router, data started flowing normally again.

3. While investigating the line 162 socket feed problems, Scott Stripling noticed a lack of data over Mexico. This was caused by a planned outage for network maintenance by Mexico which temporarily shutdown data flow through FTP inbound line 315.

HSU:

1. Robbie Berg noticed that multiple erroneous fixes from a previous EP92 storm were in the ATCF fix file (f-deck) for the new EP92. Craig Mattocks edited the file ${ATCFSTRMS}/fep922016.dat to remove the old fixes (from September), then copied the new version of the file into the ${ATCFARCHIVES}/2016 directory. Lixion Avila confirmed that the fixes imported correctly. Mike Brennan reminded us that extra fixes can also be deleted by anyone using the ATCF GUI through the "Edit Fixes" dialog box.

2. The top right monitor on the Ridges N-AWIPS workstation at the Atlantic desk went out yesterday, with the error message: “The current input timing is not supported by the monitor display. Please change your input timing to 1920x1080 @ 60Hz or any other listed timing as per the monitor specs.” An NCO Help Desk ticket has been filed.

Other:

1. NCO Help Desk tickets are not being sent out from the TSB-Admin Google Forms interface, so please submit these directly by email to:

ncwcp.helpdesk at noaa.gov

until the problem can be fixed next week.




More information about the Ncep.nhc.nco_contacts mailing list